OpenStudy (anonymous):

Write the polar equation in rectangular form. theta = π/3 A. y = 3x B. x = 3y C. y = (√3)x D. y = √(3x)

OpenStudy (anonymous):

well. my way of figuring this out which may not be right but its as close as i can do. when theta is constant the equation is a line. so the line goes in the direction of pi/3. slope is rise over run and so y/x = sin (theta)/cos(theta). sin of pi/3 is (root3)/2 and pi/3 of cosine is 1/2 so put ((root3)/2)/(1/2) = root3. so that is the slope. y=mx+b but it goes through the origin so y=(root3)x

OpenStudy (anonymous):

thanks!!!